OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2? A: y = 2x + 1 B: y = 2x + 3 C: y = 2x − 8 D: y = 2x − 3

OpenStudy (anonymous):

Plug in point (2,5). 2=x and 5=y, and determine which answer is true.

OpenStudy (anonymous):

Use the point-slope formula: \[y-y _{1}=m(x-x _{1})\] where "m" is the given slope and "x1, y1" is the given point.

OpenStudy (anonymous):

The answer will be A. 5=2*2+1 5=4+1 5=5

OpenStudy (anonymous):

Thank you! im good with slope but i was confused with this question but i get in now thank you soo much!(: